I need help with this problem...(mainly the "give three solutions part")
Solve the system.If the system has an infinite solution set, specify it and give three solutions.If the system has no solution,say so.
x+y=4(y+2)
x-y=2(y+4)
This is what I've tried(don't know if it's anywhere close to being right)...
x+y=4y+8
x-y=2y+8
x-3y=8
x-3y=8
-x+3y=-8
x-3y-x+3y=0
Then I got stuck...I think there are infinite solutions. But it says to list three of them and I don't know.
Thanks so much for your help!

Answer by unlockmath(1688)

Hello,
Let's start by getting all the x's and y's on one side of the equal sign.
x+y=4y+8
x-y=2y+8
These will look like this:
x-3y=8
x-3y=8
Since these are the exact same equations plug in a number for x and solve y.
You could then plug in another number for x and solve y. You could do this forever if you're bored or have a lot of time on your hands.
RJ
Check out a new book I wrote at:
www.math-unlock.com